OpenStudy (anonymous):

GUYS CAN YOU HELP ME ON THIS TO.ITS A PROBLEM SOLVE WITH WORDS . JEREMY IS BUILDING A LARGE DECK FOR A COMMUNITY CENTER .THE DECK IS SHAPED AS A RACTANGLE .THE WIDTH OF THE DECK IS 29 FEET .THE PERIMETER OF THE DECK IS TO BE AT LEAST 134 FEET. (A) WRITE AN INEQUALITY THAT REPRESENTS ALL POSSIBLE VALUES FOR THE LENGHT OF THE DECK . (B) FIND ALL POSSIBLE VALUES FOR THE LENGHT OF THE DECK.

OpenStudy (anonymous):

The perimeter of the deck is P = 2(L + W) = 2(L + 29), so what's the inequality that says that the perimeter is at least 134 feet?

OpenStudy (anonymous):

YES CORRECT

OpenStudy (anonymous):

134 FEET

OpenStudy (anonymous):

I BELIEVE WE HAVE TO FIND THE LENGHT ALSO

OpenStudy (anonymous):

First, take off caps lock when you type. Second, I asked a question: If I wanted to write an inequality that says "x is at least 10" then I would write \[ x \geq 10\] What would I write if I wanted to say that the perimeter is at least 134 feet?

OpenStudy (anonymous):

sorry body ok so that's my answer with (A) (B)

OpenStudy (anonymous):

Please answer my question.

OpenStudy (anonymous):

I would write that I take 134 ft divide it by 29 ft which is the width ,you will get 4.62 so the possible value is of length of the deck is 4.62 ft

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

No, you need to write an inequality like I did above.

OpenStudy (anonymous):

mhm ok lets see

OpenStudy (anonymous):

i believe p=134 (l+w)=134 (l+29

OpenStudy (anonymous):

is that correct body

OpenStudy (anonymous):

you should have >=, not =. >= means greater than or equal to, while = just means strictly equal to.

OpenStudy (anonymous):

ho ok thanks .that's all?

OpenStudy (anonymous):

That's the first part. For the second part, you solve for L. You will get L >= (something)

OpenStudy (anonymous):

and all the possible values of the length? are those

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok what do you mean about something , do i have to multiply or divide

OpenStudy (anonymous):

so the L>=which is 29 i multiply by something that gives me a perimeter p>= 134

OpenStudy (anonymous):

Okay, look. You get \[ 2(L + 29) >= 134\] so solve this inequality for L and you will get your answer.

OpenStudy (anonymous):

thank you thank you jemurray 3 ,my problem is i cant get it to put it in formulas now its easy to understand